some ordered pairs for a linear function of x are given in the table below.

x y
1 1
3 7
5 13
7 19

sorry im not good at tables

which of the following equations was used too generaTE the table above

a y=2x+1
b y=2x-1
c y=3x-2
d y=4x-3
Answer #1
you plug/ substitute x and y in each equation
start with A first
x=1 y=1 1=(2x1)+1= 3 so not same
x=1 y=7 7=(2x3)-1=5
x=5 y=13 13=(3x5)-13=2
x=7 y=19 19=(4x7)-3=19 it is the same because y=19 and then the answer you got when you plug in 7 is 19

now A equation
x=1 y=1 1=(2x1)+1=3 not the same
x=3 y=7 7=(2x3)+1=7 same

C
x=5 y=13 c y=13, x= 5 (3x5)-2=13 same

so what is left over is A with x=1 y=1
answered by: Narva